# Inference for the stochastic FitzHugh-Nagumo model from real action potential data via approximate Bayesian computation

### v1

@@ Abstract
The stochastic FitzHugh-Nagumo (FHN) model considered here is a two-dimensional nonlinear stochastic differential equation with additive degenerate noise, whose first component, the only one observed, describes the membrane voltage evolution of a single neuron. Due to its low-dimensionality, its analytical and numerical tractability, and its neuronal interpretation, it has been used as a case study to test the performance of different statistical methods in estimating the underlying model parameters. Existing methods, however, often require complete observations, non-degeneracy of the noise or a complex architecture (e.g., to estimate the transition density of the process, “recovering” the unobserved second component), and they may not (satisfactorily) estimate all model parameters simultaneously. Moreover, these studies lack real data applications for the stochastic FHN model. Here, we tackle all challenges (non-globally Lipschitz drift, non-explicit solution, lack of available transition density, degeneracy of the noise, and partial observations) via an intuitive and easy-to-implement sequential Monte Carlo approximate Bayesian computation algorithm. The proposed method relies on a recent computationally efficient and structure-preserving numerical splitting scheme for synthetic data generation, and on summary statistics exploiting the structural properties of the process. We succeed in estimating all model parameters from simulated data and, more remarkably, real action potential data of rats. The presented novel real-data fit may broaden the scope and credibility of this classic and widely used neuronal model.
@@
FitzHugh-Nagumo model, Action potential data, Hypoellipticity, Splitting numerical methods, Bayesian inference, Sequential Monte Carlo approximate Bayesian computation.
@@ 1 Introduction
In the frequentist framework, several estimation methods applicable to the stochastic FHN model have been proposed. They use adapted numerical schemes, and are based on estimation contrasts or optimal control theory. Among others, we can cite [25, 39, 47] in the case of complete observations, and [7, 9, 17, 26, 52] in the case of partial observations. In the Bayesian setting, we refer to [18, 19], where specific filtering algorithms, based on either the Taylor 1.5 scheme or the Euler-Maruyama scheme, are applied to derive the marginal posterior distributions of the FHN model parameters.
@@
As ABC relies on running extensive simulations from the model under different parameter values (step two), the quality of its results depends strongly on the choice of the numerical scheme used for synthetic data generation. Due to its simple implementation and properties (e.g. Gaussian transition densities), the standard Euler-Maruyama scheme has been used for data-simulation from any SDE model (see, e.g., [43, 44, 45, 54]), including the stochastic FHN. However, in [6] it has been recently shown that such scheme may lead to drastic fails in the estimation, due to its non-structure-preserving nature. Instead, adopting a structure-preserving numerical scheme led to successful inference of the underlying parameters (from partially observed and hypoelliptic Hamiltonian-type SDEs, including oscillatory neuronal models). Following this, here we use a numerical splitting scheme for the data-simulation within ABC. This method was recently proposed in [5], and shown to be both mean-square convergent (of order $1$) and structure-preserving, i.e., it retains the hypoellipticity, geometric ergodicity and oscillatory dynamics (amplitudes, frequencies and phases) of the stochastic FHN model, even for large time-discretisation steps. Hence, it is not only reliable, but also computationally efficient, which is particularly important when used within a simulation-based algorithm requiring millions of model simulations.
@@
Besides the chosen numerical scheme, the performance of ABC also depends heavily on the choice of suitable summary statistics for the investigated data. When applying ABC to SDE models, it is particularly important to choose summary statistics that are sensitive to small parameter changes, and robust to the intrinsic stochasticity of the model. Inspired by [6], we propose to use statistics that are invariant for repeated simulations under the same parameter setting, namely the invariant density and spectral density, which are both based on the geometric ergodicity property of the stochastic FHN model. Moreover, we show how such choice leads to more accurate approximate posterior densities than those obtained using standard canonical summaries (e.g., mean, variance and auto-correlation).
@@
Differently from [6], we do not embed here the structure-based summaries and structure-preserving numerical splitting scheme within a basic acceptance-rejection ABC algorithm. Indeed, such algorithm is very computationally wasteful, since parameter proposals are sampled from the prior distribution, which typically covers regions far away from the desired posterior distribution mass. Thus, many parameter candidates are rejected, after having spent computational time and effort to simulate the corresponding synthetic data, the summaries, and the distance measure. To improve this, several advanced ABC algorithms have been introduced over the years, such as sequential Monte Carlo (SMC) ABC (the gold standard state-of-art algorithm), Markov Chain Monte Carlo (MCMC) ABC, or sequential importance sampling (SIS) ABC, see [53] for a review. While such sequential ABC approaches have been applied to SDE models (see, e.g., [28, 43]), these works do typically not consider structure-based summaries and structure-preserving simulation methods. In this paper, we employ the structure-based ingredients within an efficient SMC-ABC scheme, yielding what we call structure-based and preserving (SBP) SMC-ABC method.
@@
Our SBP SMC-ABC method succeeds in estimating all parameters of the stochastic FHN model from simulated data, even for a relatively small number of observations and a short observation time horizon. More interestingly, we also apply our method to real data consisting of membrane voltages recorded in an adult female Sprague Dawley rat [40]. We analyse four recordings, two obtained under a resting condition, and the other two under a stimulation setting. We show the coherence of the estimated posterior distributions with respect to these experimental setups. The fitted models are then validated through goodness-of-fit plots obtained by simulations under the estimated model.
@@
While there exist only very few results fitting (or using) the FHN model to (for) real data applications, see e.g. [11] for a study on neural spike time data, and [32] for an investigation of vibration bearing data, this is, to the best of our knowledge, the first work on the considered stochastic FHN model investigating and fitting some real action potential data. This is a novel result which potentially broadens the models’ scope and credibility. In addition, note that the proposed method is not limited to the stochastic FHN model, but it can be applied to any other ergodic SDE for which a computationally tractable structure-preserving numerical method is available.
@@
The paper is organised as follows. In Section 2, we introduce the stochastic FHN model, and discuss its mathematical characteristics and structural properties. In Section 3, we introduce the SBP SMC-ABC method, detailing the algorithm, and all required key ingredients. In Section 4, we use this method to estimate the parameters of the stochastic FHN model from simulated data, under different observation settings. Section 5 details the results obtained on the real action potential data. Conclusions are reported in Section 6 and details on the algorithm are presented in the appendix.
@@ 2 Stochastic FHN model
The stochastic FHN model [5, 31] is a well-established prototype model of a single neuron (nerve cell), and dates back to the work of FitzHugh (1961) [16] and Nagumo (1962) [42]. It describes the generation of spikes (also called action potentials) at the intracellular level, and it can be obtained as solution to the following two-dimensional nonlinear SDE
@@
where $(W(t))_{t\in[0,T]}$ is a standard Wiener process, and the initial value $X_{0}=(V_{0},U_{0})^{\top}$ is a $\mathbb{R}^{2}$-valued random variable with bounded moments that is independent of $(W(t))_{t\in[0,T]}$. This is a common formulation of the model, where no noise enters directly into the $V$-component of SDE (1), and only the $U$-component is directly perturbed by it. The considered stochastic FHN model contains four parameters
@@
where $\beta,\gamma>0$ describe the position and duration of an action potential, respectively, $\epsilon>0$ is a time scale separation parameter, and $\sigma>0$ defines the noise intensity. The $V$-component of system (1) models the evolution of the membrane voltage of the neuron, which can be measured with suitable recordings techniques (cf. Section 5), while the $U$-component is an unobserved recovery variable modelling the ion channel kinetics.
@@
The drift $F(X(t))$ of the FHN model (1) is not globally Lipschitz continuous. However, as it satisfies a one-sided Lipschitz condition [4, 5], the SDE has a unique strong (unknown) solution $X(t)=(V(t),U(t))^{\top}$, $t\in[0,T]$, which is regular in the sense of [29], that is, it is defined on the whole time interval $[0,T]$, and paths of the process do not blow up to infinity in finite time. Moreover, such solution is a Markov process with (unknown) transition probability
@@ 3.1 SMC-ABC algorithm
Let $\pi(\theta)$ denote the prior density of an unknown parameter $\theta$, $p(y|\theta)$ the (unavailable) likelihood function of observations $y$, and $\pi(\theta|y)\propto p(y|\theta)\pi(\theta)$ the desired posterior density. We now recall the basic acceptance-rejection ABC algorithm presented in the introduction, to introduce the relevant notation. Such scheme consists of three steps: (a) sample $\theta^{\prime}$ from the prior $\pi(\theta)$; (b) conditioned on $\theta^{\prime}$, simulate a synthetic dataset $\tilde{y}_{\theta^{\prime}}$ from the model; (c) keep the sampled value $\theta^{\prime}$ if the distance $d(\cdot,\cdot)$ between the vector of summary statistics $s(\cdot)$ of the observed and simulated data is smaller than a user-defined tolerance level $\delta$. This leads to an approximate posterior given by
@@
and then normalised via $\omega_{r}^{(j)}=\tilde{\omega}_{r}^{(j)}/\sum_{l=1}^{N}\tilde{\omega}_{r}^{(l)}$. Here, $k_{r}(a|b)$ denotes the density of the proposal sampler $K_{r}(\cdot|b)$ evaluated in $a$.
@@
19:              Conditioned on $\theta_{j}^{*}$, simulate a synthetic dataset $\tilde{y}_{\theta_{j}^{*}}$ from the model,
@@
20:              Compute the summaries $s(\tilde{y}_{\theta_{j}^{*}})$.
21:              Calculate the distance $D_{j}=d\bigl{(}s(y),s(\tilde{y}_{\theta_{j}^{*}})\bigr{)}$.
22:         until $D_{j}<\delta_{r}$
23:         Set $\theta_{r}^{(j)}=\theta_{j}^{*}$.
24:         Set $\tilde{w}_{r}^{(j)}=\pi\left(\theta_{r}^{(j)}\right)/\sum\limits_{l=1}^{{N}}w_{r-1}^{(l)}k_{r}\left(\theta_{r}^{(j)}\Big{|}\theta_{r-1}^{(l)}\right)$.
25:     end for
26:     Normalise the weights $w_{r}^{(j)}=\tilde{w}_{r}^{(j)}/\sum\limits_{l=1}^{{N}}\tilde{w}_{r}^{(l)}$, for $j=1,\ldots,{N}$.
27:end while
28:Return a set of weighted parameters $(\theta_{r}^{(1)},\omega_{r}^{(1)}),\ldots,(\theta_{r}^{(N)},\omega_{r}^{(N)})$.
@@ 3.2 Ingredients for the SMC-ABC algorithm
The SMC-ABC Algorithm 1 requires the choice of several key ingredients: the data simulation method, the summaries (along with a suitable distance measure), the prior distribution, the proposal sampler, the threshold levels, and the stopping criterion (here, the allowed maximum number of synthetic data simulations). Among all, the first two play a crucial role, as we would like to apply an efficient numerical scheme capable of simulating from the model while preserving its key (structural) features, as well as to choose some summary statistics based on such properties. The chosen ingredients, detailed below, guarantee this, yielding the proposed SBP SMC-ABC method.
@@ 3.2.1 Structure-preserving synthetic data simulation with a splitting scheme
A crucial step within each simulation-based statistical method is the generation of synthetic datasets from the model for a given parameter value (cf. Lines $5$ and $19$ of Algorithm 1). As exact simulation schemes are not available for the stochastic FHN model (1), and standard Itô-Taylor type methods are non-convergent and/or non-preserving, we will simulate data from SDE (1) via the numerical splitting scheme recently proposed in [5].
This method is based on the idea of re-writing the drift coefficient $F(X(t))$ of the stochastic FHN model (1) as the sum of a linear and a non-linear part, and then splitting the model into two exactly solvable subequations: a linear SDE, which shares the properties of hypoellipticity and geometric ergodicity, and a non-linear ordinary differential equation (ODE) capturing the non-globally Lipschitz polynomial term. The exact solutions of these two subequations are then derived and, in every iteration step, composed in a suitable way (here, according to the Strang approach, which considers “half-steps” of the ODE-solution). Due to the exact treatment of relevant and cleverly chosen subparts of the stochastic FHN model (1), the resulting (Strang) splitting scheme preserves its structural properties (hypoellipticity, geometric ergodicity, and oscillatory dynamics). The path generation procedure required in Algorithm 1 using this splitting scheme is briefly summarized in the following. For details regarding its construction and features, the interested reader is referred to [5].
For the model simulation, we consider an equidistant time grid $0=t_{0}<t_{1}<\ldots<t_{n-1}<t_{n}=T$, with $t_{i}=i\Delta$, $\Delta=T/n$, $i=0,\ldots,n$. We set $\widetilde{X}(t_{0}):=X_{0}$, and denote by $\widetilde{X}(t_{i})$ the splitting approximation of the solution $X(t_{i})$ of SDE (1) at time $t_{i}$. We then define the value
$\kappa:=4\gamma/\epsilon-1,$ (4)
and the matrices involved in the exact solution of the linear SDE of the splitting framework, first the matrix exponential arising from the drift
@@
$c_{22}(t)$ $=$ $\frac{\sigma^{2}e^{-t}}{2\kappa}\biggl{(}\cos(\sqrt{\kappa}t)+\sqrt{\kappa}\sin(\sqrt{\kappa}t)-\frac{4\gamma}{\epsilon}+\kappa e^{t}\biggr{)}.$
Finally, we define the function that solves exactly the ODE subequation of the splitting framework
@@
The requirement $\kappa>0$ in Algorithm 2 guarantees that the underlying stochastic harmonic oscillator (i.e. the embedded linear SDE) of the FHN model (1) entering into the splitting framework is weakly damped, yielding suitable oscillatory dynamics. In our experiments, the prior will be chosen to meet this condition, see Section 3.2.4. This requirement can be however relaxed, and we refer to [5] for the corresponding expressions for $\kappa\leq 0$.

### v2

@@ Abstract
The stochastic FitzHugh-Nagumo (FHN) model is a two-dimensional nonlinear stochastic differential equation with additive degenerate noise, whose first component, the only one observed, describes the membrane voltage evolution of a single neuron. Due to its low-dimensionality, its analytical and numerical tractability and its neuronal interpretation, it has been used as a case study to test the performance of different statistical methods in estimating the underlying model parameters. Existing methods, however, often require complete observations, non-degeneracy of the noise or a complex architecture (e.g., to estimate the transition density of the process, “recovering” the unobserved second component) and they may not (satisfactorily) estimate all model parameters simultaneously. Moreover, these studies lack real data applications for the stochastic FHN model. The proposed method tackles all challenges (non-globally Lipschitz drift, non-explicit solution, lack of available transition density, degeneracy of the noise and partial observations). It is an intuitive and easy-to-implement sequential Monte Carlo approximate Bayesian computation algorithm, which relies on a recent computationally efficient and structure-preserving numerical splitting scheme for synthetic data generation and on summary statistics exploiting the structural properties of the process. All model parameters are successfully estimated from simulated data and, more remarkably, real action potential data of rats. The presented novel real-data fit may broaden the scope and credibility of this classic and widely used neuronal model.
@@
FitzHugh-Nagumo model, Action potential data, Hypoellipticity, Splitting numerical methods, Bayesian inference, Simulation-based inference, Sequential Monte Carlo approximate Bayesian computation.
@@ 1 Introduction
In the frequentist framework, several estimation methods applicable to the stochastic FHN model have been proposed. They use adapted numerical schemes and are based on estimation contrasts or optimal control theory. Among others, we can cite [27, 41, 50] in the case of complete observations and [7, 10, 19, 28, 56] in the case of partial observations. In the Bayesian setting, we refer to [20, 21], where specific filtering algorithms, based on either the Taylor 1.5 scheme or the Euler-Maruyama scheme, are applied to derive the marginal posterior distributions of the FHN model parameters.
@@
As ABC relies on running extensive simulations from the model under different parameter values (step two), the quality of its results depends strongly on the choice of the numerical scheme used for synthetic data generation. Due to its simple implementation and properties (e.g. Gaussian transition densities), the standard Euler-Maruyama scheme has been used for data-simulation from any SDE model (see, e.g., [46, 47, 48, 58]), including the stochastic FHN. However, in [6] it has been recently shown that such scheme may lead to drastic fails in the estimation, due to its non-structure-preserving nature. Instead, adopting a structure-preserving numerical scheme led to successful inference of the underlying parameters (from partially observed and hypoelliptic Hamiltonian-type SDEs, including oscillatory neuronal models). Following this, here we use a numerical splitting scheme for the data-simulation within ABC. This method was recently proposed in [5] and shown to be both mean-square convergent (of order $1$) and structure-preserving, i.e., it retains the hypoellipticity, geometric ergodicity and oscillatory dynamics (amplitudes, frequencies and phases) of the stochastic FHN model, even for large time-discretisation steps. Hence, it is not only reliable, but also computationally efficient, which is particularly important when used within a simulation-based algorithm requiring millions of model simulations.
@@
Besides the chosen numerical scheme, the performance of ABC also depends heavily on the choice of suitable summary statistics for the investigated data. When applying ABC to SDE models, it is particularly important to choose summary statistics that are sensitive to small parameter changes and robust to the intrinsic stochasticity of the model. Inspired by [6], we propose to use statistics that are invariant for repeated simulations under the same parameter setting, namely the invariant density and spectral density, which are both based on the geometric ergodicity property of the stochastic FHN model. Moreover, we show how such choice leads to more accurate approximate posterior densities than those obtained using standard canonical summaries (e.g., mean, variance and auto-correlation).
@@
Differently from [6], we do not embed here the structure-based summaries and structure-preserving numerical splitting scheme within a basic acceptance-rejection ABC algorithm. Indeed, such algorithm is very computationally wasteful, since parameter proposals are sampled from the prior distribution, which typically covers regions far away from the desired posterior distribution mass. Thus, many parameter candidates are rejected, after having spent computational time and effort to simulate the corresponding synthetic data, the summaries and the distance measure. To improve this, several advanced ABC algorithms have been introduced over the years, such as sequential Monte Carlo (SMC) ABC (the gold standard state-of-art algorithm), Markov Chain Monte Carlo (MCMC) ABC, or sequential importance sampling (SIS) ABC, see [57] for a review. While such sequential ABC approaches have been applied to SDE models (see, e.g., [30, 46]), these works do typically not consider structure-based summaries and structure-preserving simulation methods. In this paper, we employ the structure-based ingredients within an efficient SMC-ABC scheme, yielding what we call structure-based and preserving (SBP) SMC-ABC method.
@@
Our SBP SMC-ABC method succeeds in estimating all parameters of the stochastic FHN model from simulated data, even for a relatively small number of observations and a short observation time horizon. More interestingly, we also apply our method to real data consisting of membrane voltages recorded in an adult female Sprague Dawley rat [42]. We analyse four recordings, two obtained under a resting condition and the other two under a stimulation setting. We show the coherence of the estimated posterior distributions with respect to these experimental setups. The fitted models are then validated through goodness-of-fit plots obtained by simulations under the estimated model.
@@
While there exist only very few results fitting (or using) the FHN model to (for) real data applications, see e.g. [12] for a study on neural spike time data and [34] for an investigation of vibration bearing data, this is, to the best of our knowledge, the first work on the considered stochastic FHN model investigating and fitting some real action potential data. This is a novel result which potentially broadens the models’ scope and credibility. In addition, note that the proposed method is not limited to the stochastic FHN model, but it can be applied to any other ergodic SDE for which a computationally tractable structure-preserving numerical method is available.
@@
The paper is organised as follows. In Section 2, we introduce the stochastic FHN model and discuss its mathematical characteristics and structural properties. In Section 3, we introduce the SBP SMC-ABC method, detailing the algorithm and all required key ingredients. In Section 4, we use this method to estimate the parameters of the stochastic FHN model from simulated data, under different observation settings. Section 5 details the results obtained on the real action potential data. Conclusions are reported in Section 6 and details on the algorithm and results are presented in the appendix.
@@ 2 Stochastic FHN model
The stochastic FHN model [5, 33] is a well-established prototype model of a single neuron (nerve cell) and dates back to the work of FitzHugh (1961) [18] and Nagumo (1962) [44]. It describes the generation of spikes (also called action potentials) at the intracellular level. Noise can be included either on the membrane potential or on the conductance dynamics of the model. Here, we consider the randomness of opening and closing ion channels at the surface of the neuron membrane through noise acting on the conductance dynamics [45]. The resulting process can be obtained as solution to the following two-dimensional nonlinear SDE
@@
where $(W(t))_{t\in[0,T]}$ is a standard Wiener process and the initial value $X_{0}=(V_{0},U_{0})^{\top}$ is a $\mathbb{R}^{2}$-valued random variable with bounded moments that is independent of $(W(t))_{t\in[0,T]}$. This is a common formulation of the model, where no noise enters directly into the $V$-component of SDE (1) and only the $U$-component is directly perturbed by it. The considered stochastic FHN model contains four parameters
@@
where $\beta,\gamma>0$ describe the position and duration of an action potential, respectively, $\epsilon>0$ is a time scale separation parameter and $\sigma>0$ defines the noise intensity. The $V$-component of system (1) models the evolution of the membrane voltage of the neuron, which can be measured with suitable recording techniques (cf. Section 5), while the $U$-component is an unobserved recovery variable modelling the ion channel kinetics.
@@
The drift $F(X(t))$ of the FHN model (1) is not globally Lipschitz continuous. However, as it satisfies a one-sided Lipschitz condition [4, 5], the SDE has a unique strong (unknown) solution $X(t)=(V(t),U(t))^{\top}$, $t\in[0,T]$, which is regular in the sense of [31], that is, it is defined on the whole time interval $[0,T]$ and paths of the process do not blow up to infinity in finite time. Moreover, such solution is a Markov process with (unknown) transition probability
@@ 3.1 SMC-ABC algorithm
Let $\pi(\theta)$ denote the prior density of an unknown parameter $\theta$, $p(y|\theta)$ the (unavailable) likelihood function of observations $y$ and $\pi(\theta|y)\propto p(y|\theta)\pi(\theta)$ the desired posterior density. We now recall the basic acceptance-rejection ABC algorithm presented in the introduction, to introduce the relevant notation. Such scheme consists of three steps: (a) sample $\theta^{\prime}$ from the prior $\pi(\theta)$; (b) conditioned on $\theta^{\prime}$, simulate a synthetic dataset $\tilde{y}_{\theta^{\prime}}$ from the model; (c) keep the sampled value $\theta^{\prime}$ if the distance $d(\cdot,\cdot)$ between the vector of summary statistics $s(\cdot)$ of the observed and simulated data is smaller than a user-defined tolerance level $\delta$. This leads to an approximate posterior given by
@@
and then normalised via $\omega_{r}^{(j)}=\tilde{\omega}_{r}^{(j)}/\sum_{l=1}^{N}\tilde{\omega}_{r}^{(l)}$. Here, $k_{r}(a|b)$ denotes the density of the proposal sampler $K_{r}(\cdot|b)$ evaluated at $a$.
@@
19:              If $\pi(\theta_{j}^{*})=0$, return to Line 17.
20:              Conditioned on $\theta_{j}^{*}$, simulate a synthetic dataset $\tilde{y}_{\theta_{j}^{*}}$ from the model,
@@
21:              Compute the summaries $s(\tilde{y}_{\theta_{j}^{*}})$.
22:              Calculate the distance $D_{j}=d\bigl{(}s(y),s(\tilde{y}_{\theta_{j}^{*}})\bigr{)}$.
23:         until $D_{j}<\delta_{r}$
24:         Set $\theta_{r}^{(j)}=\theta_{j}^{*}$.
25:         Set $\tilde{w}_{r}^{(j)}=\pi\left(\theta_{r}^{(j)}\right)/\sum\limits_{l=1}^{{N}}w_{r-1}^{(l)}k_{r}\left(\theta_{r}^{(j)}\Big{|}\theta_{r-1}^{(l)}\right)$.
26:     end for
27:     Normalise the weights $w_{r}^{(j)}=\tilde{w}_{r}^{(j)}/\sum\limits_{l=1}^{{N}}\tilde{w}_{r}^{(l)}$, for $j=1,\ldots,{N}$.
28:end while
29:Return a set of weighted parameters $(\theta_{r}^{(1)},\omega_{r}^{(1)}),\ldots,(\theta_{r}^{(N)},\omega_{r}^{(N)})$.
@@ 3.2 Ingredients for the SMC-ABC algorithm
The SMC-ABC Algorithm 1 requires the choice of several key ingredients: the data simulation method, the summaries (along with a suitable distance measure), the prior distribution, the proposal sampler, the threshold levels and the stopping criterion (here, the allowed maximum number of synthetic data simulations). Among all, the first two play a crucial role, as we would like to apply an efficient numerical scheme capable of simulating from the model while preserving its key (structural) features, as well as to choose some summary statistics based on such properties. The chosen ingredients, detailed below, guarantee this, yielding the proposed SBP SMC-ABC method.
@@ 3.2.1 Structure-preserving synthetic data simulation with a splitting scheme
A crucial step within each simulation-based statistical method is the generation of synthetic datasets from the model for a given parameter value (cf. Lines $5$ and $20$ of Algorithm 1). As exact simulation schemes are not available for the stochastic FHN model (1) and standard Itô-Taylor type methods are non-convergent and/or non-preserving, we will simulate data from SDE (1) via the numerical splitting scheme recently proposed in [5].
This method is based on the idea of re-writing the drift coefficient $F(X(t))$ of the stochastic FHN model (1) as the sum of a linear and a non-linear part, and then splitting the model into two exactly solvable subequations: a linear SDE, which shares the properties of hypoellipticity and geometric ergodicity, and a non-linear ordinary differential equation (ODE) capturing the non-globally Lipschitz polynomial term. The exact solutions of these two subequations are then derived and, in every iteration step, composed in a suitable way (here, according to the Strang approach, which considers “half-steps” of the ODE-solution). Due to the exact treatment of relevant and cleverly chosen subparts of the stochastic FHN model (1), the resulting (Strang) splitting scheme preserves its structural properties (hypoellipticity, geometric ergodicity and oscillatory dynamics). The path generation procedure required in Algorithm 1 using this splitting scheme is briefly summarized in the following. For details regarding its construction and features, the interested reader is referred to [5].
For the model simulation, we consider an equidistant time grid $0=t_{0}<t_{1}<\ldots<t_{n-1}<t_{n}=T$, with $t_{i}=i\Delta$, $\Delta=T/n$, $i=0,\ldots,n$. We set $\widetilde{X}(t_{0}):=X_{0}$ and denote by $\widetilde{X}(t_{i})$ the splitting approximation of the solution $X(t_{i})$ of SDE (1) at time $t_{i}$. The considered splitting scheme relies on the linear SDE $dX_{t}=AX(t)+\Sigma dW(t)$, with
$A=\begin{pmatrix}0&-\frac{1}{\epsilon}\\ \gamma&-1\end{pmatrix},$
whose exact solution depends on the matrix exponential $E(t):=e^{At}$ arising from the drift coefficient and the covariance matrix
[…]
The resulting process constitutes a weakly-, critically-, or over-damped stochastic harmonic oscillator, depending on whether the value
$\kappa:=\frac{4\gamma}{\epsilon}-1$ (5)
is positive, zero, or negative, respectively, see e.g. Chapter 5 of [61]. Requiring $\kappa>0$ (i.e., considering the weakly damped case), guarantees that the model produces suitable oscillatory dynamics. In this case, the matrix exponential and covariance matrix are given by
@@
$c_{22}(t)$ $=$ $\frac{\sigma^{2}e^{-t}}{2\kappa}\biggl{(}\cos(\sqrt{\kappa}t)+\sqrt{\kappa}\sin(\sqrt{\kappa}t)-\frac{4\gamma}{\epsilon}+\kappa e^{t}\biggr{)},$
respectively, for $t>0$. Finally, we define the function that solves exactly the ODE subequation of the splitting framework
@@
In our experiments, the prior will be chosen to meet the condition $\kappa>0$ in Algorithm 2, see Section 3.2.4. However, this requirement can be relaxed, and we refer to [5] for the corresponding expressions of $E(t)$ and $C(t)$ for $\kappa\leq 0$.